D Max Corporation’s bond has a 7% coupon rate and a $1,000 face value. The coupon...

  1. D Max Corporation’s bond has a 7% coupon rate and a $1,000 face value. The coupon is paid semi-annually and the bond has 25 years to maturity. If the bond holders required rate of return is 6% per annum

    a.Calculate the value of the bond
    b.identify whether it is a premium, par or a discount bond.
Answer #1
Period Discounting Factor
[1/(1.03^period)]
Discounting Factor Annuity
(Sum of discounting factor & all previous discounting factors)
1 0.970873786
2 0.942595909
3 0.915141659
4 0.888487048
5 0.862608784
6 0.837484257
7 0.813091511
8 0.789409234
9 0.766416732
10 0.744093915
11 0.722421277
12 0.70137988
13 0.68095134
14 0.661117806
15 0.641861947
16 0.623166939
17 0.605016446
18 0.587394608
19 0.570286027
20 0.553675754
21 0.537549276
22 0.521892501
23 0.506691748
24 0.491933736
25 0.477605569
26 0.463694727
27 0.450189056
28 0.437076753
29 0.424346362
30 0.41198676
31 0.399987145
32 0.388337034
33 0.377026247
34 0.3660449
35 0.355383398
36 0.345032425
37 0.334982937
38 0.325226152
39 0.315753546
40 0.306556841
41 0.297628001
42 0.288959224
43 0.280542936
44 0.272371782
45 0.264438624
46 0.256736528
47 0.249258765
48 0.241998801
49 0.234950292
50 0.22810708 25.72976401

a)

Value of Bond = PV of All Coupons + PV of Maturity Amount

= (1000*3.5%)*PV Annuity Factor @3% for 50 periods + 1000*PV Discounting Factor @3% for 50 periods

= 35*25.729764 + 1000*0.228107

= 900.54174 + 228.107

= $1128.64874

b) Bond Value is HIGHER than Par Value. Therefore, Premium Bond.